### v1

@@ 1 INTRODUCTION
The Internet of Things (IoT) has emerged as one of the most important technologies of the twenty-first century. In recent years, the number of new IoT products and applications has grown exponentially, and IoT technology has been extensively applied across diverse fields such as smart cities, smart transportation, smart factories, smart healthcare, intelligent warning systems, and disaster management [1]. Automated devices such as sensors, tracking devices, and meters, often referred to as machine-type communication (MTC) devices, are assumed to have a density of one million devices per square kilometer [2, 3, 4]. These networks can be characterized by a massive amount of data available through the devices and the possibility of exchange of this information among the large number of devices without or with little intervention by human [1, 5]. Although significant progress has been made by the 3rd generation partnership project (3GPP) to make 5G mobile and wireless networks support MTC, it can still be a challenging task to transfer the large amount of data collected by massive number of devices using the existing infrastructures [3, 4, 5, 6, 7].
@@
With the popularity of MTC applications, significant research is underway to upgrade the existing algorithms that are adopted in long term evolution advanced (LTE-A) for MTC for use in 5G networks. These algorithms are categorized as either contention-based or contention-free methods. Among them, the contention-based multi-channel slotted ALOHA scheme is widely used in uplink random access channels (RACH) [8, 9, 10, 11]. This is due to the scheme’s low a priori knowledge requirement, in which devices randomly select channels and transmit preambles/packets to the associated base station without negotiation [12]. The contention-based approach is suitable for a network with limited number of access requests [13]. However, this scheme may fail in massive MTC (mMTC) networks during bursty traffic in which tens of thousands devices need to reserve the wireless access resources and transmit their packet in a short period of time [13, 6, 14, 15]. Collisions are thus the primary issue with the contention-based slotted ALOHA schemes in mMTC networks [10].
@@
On the other hand, contention-free schemes suffer from high complexity and scalability issues. For example, in grant-free schemes [16, 17, 18], a user is immediately assigned a unique non-orthogonal preamble (similar to an identification number) upon joining the network and uses this preamble to transmit its packets to the network. Employing no-orthogonal preambles resolves the collision issue in the network at the cost of increased user detection complexity at the receiver. Moreover, the total number of devices that can exist in a network is limited by the available non-orthogonal preambles [19]. In this work, we focus on designing a method to address the congestion issue in contention-based multi-channel ALOHA schemes.
@@ 1.2 Prior Works
In [23] and [51], the authors proposed an ML based method to predict the MTC network traffic during the bursty scenarios. Their work focuses exclusively on predicting traffic pattern after observing an event and without considering the uniform traffic of the devices. However, bursty and uniform traffic behave differently, and while neither alone may cause congestion, the combination of the two can. This necessitates the development of a ML network capable of predicting both uniform and bursty traffic in real-world scenarios.
@@ 1.3 Traditional Rolling Algorithm
Although LSTM model outperforms typical RNN networks, it cannot maintain its accuracy over a longer period of time in the multi-step prediction of the time-series [33]. Consequently, prediction length in offline methods are limited to a few steps based of the the time-series properties. On the other hand, longer and continuous predictions are possible in live scenarios where fresh data is available every often. In live scenarios, the traditional rolling method is often utilized for live predictions. This method involves frequent updates to historical data by concatenating new data while removing the oldest samples. Subsequently, similar to offline methods, the LSTM network is initialized with the updated historical data, enabling predictions for the next few steps. The drawbacks of this method include the need for a large buffer to store historical data, extra computational overhead due to updating the historical data, and the time-consuming nature of initializing the LSTM with extended historical data. These factors make this method undesirable for time-sensitive and resource-limited applications. Moreover, a smaller buffer size deprives the longer dependencies from the time-series sequence and thereby reduces prediction accuracy. To address these challenges, we propose a novel algorithm that replaces the traditional rolling method, offering a lighter and faster approach for live predictions. The idea behind this approach is that all the extracted information from the historical data is abstracted in hidden states of LSTM network. Thus, instead of extracting the dependencies from the updated historical data upon arrival of fresh data, we restore the LSTM states and update them with the fresh data. This makes the initialization process much faster than rolling method.
@@ 1.5 Organization
This paper is organized as follows. In section 2, the system model and definitions are specified. The ML networks for traffic pattern and congestion prediction are described in section 3. The FLSP algorithm is explained in details in section 4. Training procedure of the proposed burst detection network is described in section 5. Simulation results are presented and discussed in section 6. Finally, conclusion is provided in section 7.
@@ 2.2 Definitions
3 Propose Architecture
Fig. 3 depicts the two main networks in our proposed ML architecture. The first network comprises a two-layer LSTM and a two-layer DenseNet with dropout layers introduced between the LSTM and DenseNet sections, as well as between the connections of the DenseNet layer, to mitigate overfitting during training. The combination of DenseNet and LSTM ensures a high convergence rate and effective extraction of both short-term and long-term dependencies in the traffic pattern.
@@ 3 Propose Architecture
4 Online Mode Prediction
In this section, we provide a comprehensive description of the Fast LiveStream Predictor (FLSP) algorithm, a novel approach designed to replace traditional rolling algorithm in time-series prediction problem in live scenarios. We also provide the complexity analysis of the proposed algorithm and compare it to the rolling algorithm.
4.1 Proposed Algorithm
@@
After initialization, the LSTM network is ready to generate predictions. A recursive algorithm is employed in the forward direction to predict sequential patterns. Based on this algorithm, the network’s output from previous steps is used as input to predict the output at the current step. The recursive algorithm is depicted in Fig. 4 and is explained by the following equation:
@@
Fig. 5 illustrates an exemplary implementation of the rolling method and FLSP algorithm. In both methods, the LSTM network takes an input sequence, depicted by the solid black frame at each step. Notably, the FLSP algorithm uses significantly shorter input data for the LSTM network compared to the rolling method, resulting in reduced processing time. At each step of both methods, predictions are generated for a specific number of steps into the future. However, only the most recent steps are retained and appended to the output sequence. For elaborate, assume that the network predicts $l_{p}$ steps ahead using the recursive method at each step after feeding the actual data. Then, the system remains in a state of readiness to receive the most recent data of size $l_{f}$, collected from the mMTC network (fresh data), in a way that $l_{f}<l_{p}$. After updating the LSTM states with the fresh data, the recursive algorithm generates the predictions for the next $l_{p}$ steps while retaining only the last $l_{f}$ steps of the predicted data. The reason for keeping only the last $l_{f}$ steps is that we are primarily interested in the unseen information, specifically the time slots that have not been predicted before, so we ignore the rest of the points when generating the output sequence. It is important to note that the determination of prediction parameters depends on network configurations, the frequency of collecting fresh data, the required forecast length, and the desired prediction accuracy.
@@ 4.2 Time Complexity
5 Training the burst detection network
@@
The burst detection network is designed to predict the occurrence of bursty congestion in the following $l_{p}$ slots. This network gets output chunks of data from the traffic prediction network and determines whether the predicted traffic samples belong to a bursty event. A chunk of data is defined as the concatenation of fresh data (of size $l_{f}$) and predicted samples (of size $l_{p}$) at each step. An exemplary chunk of data is shown by a yellow dashed frame in Fig. 5(b). The chunks of data are preferred over the output sequence ($\widehat{y}$) as they are smoother and more accurate than the traffic sequence. The samples in the output sequence of the traffic prediction network ($\widehat{y}$) are at least $l_{p}-l_{f}$ and at most $l_{p}$ steps away from the last sample of the real traffic (see Fig. 5(b)). According to [33], the prediction accuracy decreases as the number of steps increases. Thus, we prefer to use the data chunks where the predicted points are at least one step and at most $l_{p}$ steps away from the most recent samples of the real traffic, providing a smaller average lead time from the real data. Moreover, generating data chunks in a single round results in a smoother pattern compared to the output sequence ($\widehat{y}$), which is composed of multiple data segments. This approach aligns with the recommendation of researchers in [23, 51] who suggest using smoother input sequences in machine learning problems to enhance learning efficiency. Simulations on the validation data show that when $T_{pred}=2\ [sec]$ ($l_{p}=400$ time slots), using chunks of data rather than predicted streams results in a $76\%$ improvement, confirming our statement.
@@ 6 Simulation Results
The simulation setup is described in this section, followed by evaluation metrics, simulation results and discussion.

### v2

@@ 1 INTRODUCTION
The Internet of Things (IoT) has emerged as one of the most important technologies of the twenty-first century. In recent years, the number of new IoT products and applications has grown exponentially, and IoT technology has been extensively applied across diverse fields such as smart cities, smart transportation, smart factories, smart healthcare, intelligent warning systems, and disaster management [1]. Automated devices such as sensors, tracking devices, and meters, often referred to as machine-type communication (MTC) devices, are assumed to have a density of one million devices per square kilometer [2, 3, 4]. These networks can be characterized by a massive amount of data available through the devices and the possibility of exchange of this information among the large number of devices without or with little intervention by human [1]. Examples of MTC application in smart manufacturing and industrial applications are discussed in [5] and [6], demonstrating significant improvements in operational efficiency through real-time data acquisition, fault detection, and intelligent decision-making. However, these advancements also result in increased traffic from numerous interconnected devices, highlighting the need for effective traffic management to prevent network congestion. Although significant progress has been made by the 3rd generation partnership project (3GPP) to make 5G mobile and wireless networks support MTC, it can still be a challenging task to transfer the large amount of data collected by massive number of devices using the existing infrastructures [7].
@@
With the popularity of MTC applications, significant research is underway to upgrade the existing algorithms that are adopted in long term evolution advanced (LTE-A) for MTC for use in 5G networks. These algorithms are categorized as either contention-based or contention-free methods. Among them, the contention-based multi-channel slotted ALOHA scheme is widely used in uplink random access channels (RACH) [8, 9]. This is due to the scheme’s low a priori knowledge requirement, in which devices randomly select channels and transmit preambles/packets to the associated base station without negotiation [10]. The contention-based approach is suitable for a network with limited number of access requests [11]. However, this scheme may fail in massive MTC (mMTC) networks during bursty traffic in which tens of thousands devices need to reserve the wireless access resources and transmit their packet in a short period of time [11]. Collisions are thus the primary issue with the contention-based slotted ALOHA schemes in mMTC networks [12].
@@
On the other hand, contention-free schemes suffer from high complexity and scalability issues. For example, in grant-free schemes [13, 14], a user is immediately assigned a unique non-orthogonal preamble (similar to an identification number) upon joining the network and uses this preamble to transmit its packets to the network. Employing no-orthogonal preambles resolves the collision issue in the network at the cost of increased user detection complexity at the receiver. Moreover, the total number of devices that can exist in a network is limited by the available non-orthogonal preambles [15]. In this work, we focus on designing a method to address the congestion issue in contention-based multi-channel ALOHA schemes.
@@ 1.2 Prior Works
In [19] and [46], the authors proposed an ML based method to predict the MTC network traffic during the bursty scenarios. Their work focuses exclusively on predicting traffic pattern after observing an event and without considering the uniform traffic of the devices. However, bursty and uniform traffic behave differently, and while neither alone may cause congestion, the combination of the two can. This necessitates the development of a ML network capable of predicting both uniform and bursty traffic in real-world scenarios. Table 1 summarizes the related works on traffic prediction in MTC networks.
Table 1: Summary of Related Work on Traffic Prediction in mMTC Networks.
Reference Main Focus Prediction Model Target Traffic Type MTC Groups Online Prediction
[19] Traffic prediction LSTM Aggregated traffic Single ×
[27]
Traffic prediction,
resource allocation
Direct information Source traffic Single ×
[36] Resource allocation
Statistical models
(fluid model extension,
crowd-source information,
recursive estimation)
Aggregated traffic Single ×
[39]
Traffic prediction,
resource allocation
LSTM Source traffic Single ×
[40] Resource allocation LSTM, support vector machine (SVM) Source traffic Single ×
[41] Resource allocation Statistical models Source traffic Single ×
[42] Resource allocation On-Off Markov process Source traffic Single ×
[43] Resource allocation LSTM, SVM, KNN Aggregated traffic Single ×
[46] Traffic prediction LSTM Aggregated traffic Single ×
This work Traffic prediction LSTM, GRU Aggregated traffic Multiple
@@ 1.3 Traditional Rolling Algorithm
Although LSTM model outperforms typical recurrent neural networks (RNNs), it cannot maintain its accuracy over a longer period of time in the multi-step prediction of the time-series [29]. Consequently, prediction length in offline methods are limited to a few steps based of the the time-series properties. On the other hand, longer and continuous predictions are possible in live scenarios where fresh data is available every often. In live scenarios, the traditional rolling method is often utilized for live predictions. This method involves frequent updates to historical data by concatenating new data while removing the oldest samples. Subsequently, similar to offline methods, the LSTM network is initialized with the updated historical data, enabling predictions for the next few steps. The drawbacks of this method include the need for a large buffer to store historical data, extra computational overhead due to updating the historical data, and the time-consuming nature of initializing the LSTM with extended historical data. These factors make this method undesirable for time-sensitive and resource-limited applications. Moreover, a smaller buffer size deprives the longer dependencies from the time-series sequence and thereby reduces prediction accuracy. To address these challenges, we propose a novel algorithm that replaces the traditional rolling method, offering a lighter and faster approach for live predictions. The idea behind this approach is that all the extracted information from the historical data is abstracted in hidden states of LSTM network. Thus, instead of extracting the dependencies from the updated historical data upon arrival of fresh data, we restore the LSTM states and update them with the fresh data. This makes the initialization process much faster than rolling method.
@@ 1.4 Contributions
1.5 Acronyms
To improve readability, all acronyms used in this paper are defined in Table 2. The acronyms are sorted in alphabetical order for ease of reference.
Table 2: List of Acronyms (sorted alphabetically)
Acronym Definition
3GPP 3rd generation partnership project
5G Fifth generation of cellular technology
ACB Access class barring
AOI Age of information
CNN-1D One-dimensional convolutional neural network
EAB Extended access barring
FFNN Feed-forward neural network
FLOP Floating-point operation
FLSP Fast LiveStream Predictor
GRU Gated recurrent unit
IoT Internet of Things
LSTM Long-term short-term memory
LTE-A Long term evolution advanced
ML Machine learning
mMTC Massive machine-type communication
MSE Mean square error
MTC Machine-type communication
NOMA Non-orthogonal multiple access
NR New radio
OFDMA Orthogonal frequency division multiple access
RACH Random access channels
RNN Recurrent neural network
@@ 1.6 Organization
This paper is organized as follows. In section 2, the system model and definitions are specified. The ML networks for traffic pattern and congestion prediction are described in section 3. The FLSP algorithm is explained in details in section 3.2. Training procedure of the proposed burst detection network is described in section 3.5. Simulation results are presented and discussed in section 4. Finally, conclusion is provided in section 5.
@@ 2.2 Definitions
3 Proposed Architecture
In this section, we explain the proposed ML architecture and provide a comprehensive description of the Fast LiveStream Predictor (FLSP) algorithm, a novel approach designed to replace traditional rolling algorithm in time-series prediction problem in live scenarios. We also provide the complexity analysis of the proposed algorithm and compare it to the rolling algorithm.
3.1 Proposed ML Architecture
Figure. 3 depicts the two main networks in our proposed ML architecture. The first network comprises a two-layer LSTM and a two-layer DenseNet with dropout layers introduced between the LSTM and DenseNet sections, as well as between the connections of the DenseNet layer, to mitigate overfitting during training. The combination of DenseNet and LSTM ensures a high convergence rate and effective extraction of both short-term and long-term dependencies in the traffic pattern.
@@ 3.1 Proposed ML Architecture
3.2 Proposed Algorithm for Online Mode Prediction
@@
After initialization, the LSTM network is ready to generate predictions. A recursive algorithm is employed in the forward direction to predict sequential patterns. Based on this algorithm, the network’s output from previous steps is used as input to predict the output at the current step. The recursive algorithm is depicted in Figure. 4 and is explained by the following equation:
@@
Fig. 5 illustrates an exemplary implementation of the rolling method and FLSP algorithm. In both methods, the LSTM network takes an input sequence, depicted by the solid black frame at each step. Notably, the FLSP algorithm uses significantly shorter input data for the LSTM network compared to the rolling method, resulting in reduced processing time. At each step of both methods, predictions are generated for a specific number of steps into the future. However, only the most recent steps are retained and appended to the output sequence. Assume that the network predicts $l_{p}$ steps ahead using the recursive method at each step after feeding the actual data. Then, the system remains in a state of readiness to receive the most recent data of size $l_{f}$, collected from the mMTC network (fresh data), in a way that $l_{f}<l_{p}$. After updating the LSTM states with the fresh data, the recursive algorithm generates the predictions for the next $l_{p}$ steps while retaining only the last $l_{f}$ steps of the predicted data. The reason for keeping only the last $l_{f}$ steps is that we are primarily interested in the unseen information, specifically the time slots that have not been predicted before, so we ignore the rest of the points when generating the output sequence. For example, when $l_{p}=200$ ($T_{pred}=1[sec]$), $l_{f}=100$ ($T_{fresh}=0.5[sec]$), and $l_{buff}=200$, the FLSP algorithm only receives fresh data of size $l_{f}=100$ time slots as input to the prediction model. In contrast, the rolling algorithm requires $l_{buff}+l_{f}=300$ time slots of data as input, which includes both buffered and fresh data. Both algorithms generate output for $l_{p}=200$ time slots, but only the last $l_{f}=100$ time slots are retained and appended to the output sequence. This results in an output sequence that aligns with the target sequenced used to calculate the evaluation metrics. The target sequence is the shifted version of the input sequence by $l_{p}$ time slots, representing the expected output from the prediction model at each time (ground truth data). It is important to note that the determination of prediction parameters depends on network configurations, the frequency of collecting fresh data, the required forecast length, and the desired prediction accuracy.
@@ 3.3 Time Complexity
3.4 Floating-point Operations and Number of Parameters
In this section, we provide the floating point operations (FLOPs) required by the FLSP and the rolling algorithm to generate the output data. For comparative analysis, we also calculate the FLOPS required by two additional models, the gated recurrent unit (GRU) and the One-dimensional convolutional neural network (CNN-1D). Both GRU and LSTM models, as variants of RNNs, are adept at processing sequential data. However, the GRU architecture is simpler, featuring only two gates. CNN-1D models, applicable to time-series prediction problems, focus primarily on extracting local dependencies within the input data. Unlike RNN-based models that capture broader temporal patterns, CNN-1D models excel in identifying short-term correlations. A limitation of CNN-1D models is that they are only compatible with the rolling algorithm. Additionally, the number of parameters for each model is provided, reflecting their training complexity. The FLOPs required by the LSTM-DenseNet network when employing the rolling and FLSP algorithms are detailed as follows:
$\begin{split}FLOP_{Roll}^{(LSTM)}=&\frac{l_{buff}+l_{p}}{l_{f}}*(\sum_{l=1}^{L_{LSTM}}(8h_{l}(h_{l-1}+h_{l})\\ &+29h_{l})+\sum_{l=1}^{L_{D}}(2m_{l}m_{l-1})),\end{split}$ (10)
$\begin{split}FLOP_{FLSP}^{(LSTM)}=&\frac{l_{f}+l_{p}}{l_{f}}*(\sum_{l=1}^{L_{LSTM}}(8h_{l}(h_{l-1}+h_{l})\\ &+29h_{l})+\sum_{l=1}^{L_{D}}(2m_{l}m_{l-1})),\end{split}$ (11)
and the number of the model’s parameters is:
$\begin{split}P_{LSTM}=\sum_{l=1}^{L_{LSTM}}4h_{l}(h_{l-1}+h_{l}+2)+\sum_{l=1}^{L_{D}}(m_{l}(m_{l-1}+1))\end{split}$ (12)
where, $h_{l}$ represents the hidden size of layer $l$, $h_{0}$ is the input feature size, $L_{LSTM}$ denotes the number of LSTM layers, $L_{D}$ indicates the number of linear layers in the DenseNet network, $m_{l-1}$ and $m_{l}$ respectively correspond to the input and output data sizes of layer $l$ in the DenseNet network. Following the methodology used for the LSTM-DenseNet model, we derive the FLOP calculations and parameter count for the GRU-DenseNet model as follows:
$\begin{split}FLOP_{Roll}^{(GRU)}=&\frac{l_{buff}+l_{p}}{l_{f}}*(\sum_{l=1}^{L_{GRU}}(6h_{l}(h_{l-1}+h_{l})+22h)\\ &+\sum_{l=1}^{L_{D}}(2m_{l}m_{l-1})),\end{split}$ (13)
$\begin{split}FLOP_{FLSP}^{(GRU)}=&\frac{l_{f}+l_{p}}{l_{f}}*(\sum_{l=1}^{L_{GRU}}(6h_{l}(h_{l-1}+h_{l})+22h)\\ &+\sum_{l=1}^{L_{D}}(2m_{l}m_{l-1})),\end{split}$ (14)
$P_{GRU}=\sum_{l=1}^{L_{GRU}}3h_{l}(h_{l-1}+h_{l}+2)+\sum_{l=1}^{L_{D}}(m_{l}(m_{l-1}+1)),$ (15)
where $L_{GRU}$ indicates the number of GRU layers. The FLOP calculations and parameter count for the CNN-1D model can be expressed as follows:
$\begin{split}FLOP&{}_{Roll}^{(CNN-1D)}=\frac{1}{l_{f}}*(\sum_{l=1}^{L_{CNN-1D}}(\frac{W}{2^{l}}C_{l}(2C_{l-1}K_{l}\\ &+1/2))+\sum_{l=1}^{L_{FC}}(2m_{l}m_{l-1}+m_{l})-m_{L_{FC}}),\end{split}$ (16)
$\begin{split}P_{CNN-1D}=&\sum_{l=1}^{L_{CNN-1D}}(C_{l}(C_{l-1}K+1))\\ &+\sum_{l=1}^{L_{FC}}(m_{l}(m_{l-1}+1)),\end{split}$ (17)
where $L_{CNN-1D}$ represents the number of CNN-1D cells, $W$ is the input data size (window size), $C_{l-1}$ and $C_{l}$ are, respectively, the input and output feature sizes of each CNN-1D cell, $K_{l}$ denotes the kernel size, $L_{FC}$ indicates the number of fully connected layers, and $m_{l-1}$ and $m_{l}$ are, respectively, the input and output data sizes of the layer $l$ in the fully connected network.
3.5 Training the Burst Detection Network
@@
The burst detection network is designed to predict the occurrence of bursty congestion in the following $l_{p}$ slots. This network gets output chunks of data from the traffic prediction network and determines whether the predicted traffic samples belong to a bursty event. A chunk of data is defined as the concatenation of fresh data (of size $l_{f}$) and predicted samples (of size $l_{p}$) at each step. An exemplary chunk of data is shown by a yellow dashed frame in Fig. 5(b). The chunks of data are preferred over the output sequence ($\widehat{y}$) as they are smoother and more accurate than the traffic sequence. The samples in the output sequence of the traffic prediction network ($\widehat{y}$) are at least $l_{p}-l_{f}$ and at most $l_{p}$ steps away from the last sample of the real traffic (see Fig. 5(b)). Therefor, the input to the burst detection network has a feature size of $ch\_size=2\times(l_{f}+l_{p})$, where the factor of $2$ accounts for the two feature in the output of the traffic prediction network: successfully detected preambles and congested preambles. According to [29], the prediction accuracy decreases as the number of steps increases. Thus, we prefer to use the data chunks where the predicted points are at least one step and at most $l_{p}$ steps away from the most recent samples of the real traffic, providing a smaller average lead time from the real data. Moreover, generating data chunks in a single round results in a smoother pattern compared to the output sequence ($\widehat{y}$), which is composed of multiple data segments. This approach aligns with the recommendation of researchers in [19, 46] who suggest using smoother input sequences in machine learning problems to enhance learning efficiency. Simulations on the validation data show that when $T_{pred}=2\ [sec]$ ($l_{p}=400$ time slots), using chunks of data rather than predicted streams results in a $76\%$ improvement, confirming our statement.
@@ 4 Simulation Results
The simulation setup is described in this section, followed by evaluation metrics, simulation results and discussion. The simulation codes used in this study are available at Github [^1] https://github.com/HosseinMehriB/mMTC-Traffic-Prediction for reproducibility and further research.
@@ 4.1 Simulation Setup
4.2 Data Labeling for Burst Detection Network
To label the data for the burst detection network, we adopted the definition provided in Subsection 2.2.2, where an area is labeled as congested if the number of detected attempts is lower than the number of arrivals within a specified time interval. In the simulations, the average number of collided attempts (calculated as the difference between total attempts and successfully detected preambles) was measured over a sliding window of approximately $3$ seconds. This value was then compared to an experimentally determined threshold of $7000$ to identify congested areas. The labeling process, based on these experimental values, demonstrated high accuracy in identifying congested regions. Then, for each prediction time $T_{pred}$, the model is expected to detect congestion $T_{pred}$ seconds prior to its start. Accordingly, the expected congestion labels for the given $T_{pred}$ were assigned a value of “$1$” if congestion occurred within the next $T_{pred}$ seconds, and “00” otherwise. Fig. 6 shows an example of the actual congested area and its corresponding expected congestion labels for $T_{pred}=2$ seconds.
Figure 6: An example of labeling the congested area for $T_{pred}=2$ seconds. The extracted congestion labels (green dashed line) accurately delineate the congested region where there is a gap between the arrivals and successfully detected preambles.